OpenStudy (anonymous):

Use the sum formula to find the exact value of sin(aplha+beta) under the given conditions.

OpenStudy (anonymous):

\[\sin \alpha=4/5, 0<\alpha<\pi/2, \cos \beta=(6\sqrt{85})/85, -\pi/2<\beta<0\]

OpenStudy (anonymous):

\[\sin (\alpha+\beta)=(3\sqrt{85})/425\]

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

ya corret